how to derive the error equation for F = mg. BACK SUBSTITUTE & simplify.
Answer #1

Solution ) we have momentum P =mv

On differentiating with respect to time 't'

dP/dt = d(mv)/dt = m(dv/dt) = ma

dv/dt = acceleration a

we have un Cerfainity (o means ero r mau plus mass Tmes on MaJA the e